OpenStudy (anonymous):

What is √180 rounded to the nearest hundredth?Select one of the options below as your answer: A. 13.41 B. 13.42 C. 13.43 D. 13.44

OpenStudy (hlambach):

put it into your calculator

OpenStudy (hlambach):

or do you not know how?

OpenStudy (hlambach):

if your calculator is like the one below, press 2nd and x^2

OpenStudy (hlambach):

and if you cannot use your calculator: √180 Simplify the radical √180 = √(36 x 5) Since 36 is a perfect square 6√5 The square root of five is 2.236... plus a ton of other numbers So 6 x 2.23 = 13.416... And rounded to the nearest tenth 13.42

OpenStudy (hlambach):

Hope that helped!

OpenStudy (amistre64):

another method is to make a guess \[\frac{180+g^2}{2g}\] let g=13 \[\frac{180+13^2}{2*13}=new~guess\] run the new guess thru .. after 2 times you should be able to see it limiting to the correct answer